OpenStudy (anonymous):

Converting a equation of a circle in standard form to find center points and radius: Okay so the equation is x^2-7x+y^2+10y+16=0. I already solved the left side of the equation (x-7/2)^2+(y+5)^2 so my center is (7/2, -5) but once I moved the 16 to the right so that it's =-16 am not quite sure what to do now .... ? Help am so close.

OpenStudy (anonymous):

you missed while completing the square note: the number 16 will add to other quantities so the final equation has not 16

OpenStudy (anonymous):

Am not quite sure what you mean.... and I know the final product isn't 16 but that's why am not sure what to do with it :( @esamalaa

OpenStudy (anonymous):

you can write your solution steps and i will find errors for you

OpenStudy (anonymous):

\[X ^{2}-7x+Y ^{2}+10y+16=0\]

OpenStudy (anonymous):

ok what is the first step ?

OpenStudy (anonymous):

Sorry its taking me forever. .... my laptop keeps on stalling and I just moved to my phone..... @esamalaa

OpenStudy (anonymous):

X^2 - 7X + y^2 +10y=-16

OpenStudy (anonymous):

X^2-7x+(7/2)^2. +y^2+10y+(10/2)^2

OpenStudy (anonymous):

(X-7/2)^2 + (y+5)^2 = -16

OpenStudy (anonymous):

Center (7/2, -5)

OpenStudy (anonymous):

=-16 +7/2 -5 <<<<< is this what am suppose to do next?

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

-16 +(7/2)^2+25

OpenStudy (anonymous):

Ohhh is it 85/4 ??? @esamalaa

OpenStudy (anonymous):

yes but that is the radius squared if you want the radius itself take the sq root for 85/4

OpenStudy (anonymous):

Yay ! So if I did it the other way would it be 2.304..... ? @esamalaa

OpenStudy (anonymous):

4.60

OpenStudy (anonymous):

Oh ok I see ! Yay ! Thank you so much, am new to this fourm and you have been a big help ! God Bless You ! @esamalaa

OpenStudy (anonymous):

anytime